OpenStudy (anonymous):

∫∫(-xy^3 + 6y^2) dA where R = {(x,y)|0<=x <=3, 0<= y <=1} calculate the double integral.

OpenStudy (sirm3d):

\[\large \int\limits_{y=0}^{y=1}\int\limits_{x=0}^{x=3}(-xy^3+6y^2)dxdy\]or \[\large \int\limits_{0}^{3}\int\limits_{0}^{1}(-xy^3+6y^2)dydx\]

OpenStudy (anonymous):

i worked out. is the ans -33/8.

OpenStudy (anonymous):

15/8

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

i have solved

OpenStudy (anonymous):

is it possible to show ur working. when recalculated i got 39/8.

OpenStudy (sirm3d):

39/8 it is.

OpenStudy (anonymous):

yes 39/8

OpenStudy (anonymous):

ok . thanks